OpenStudy (anonymous):

A coyote chasing a rabbit is moving 8.00 m/s due east at one moment and 8.80 m/s due south 4.20s later.

OpenStudy (anonymous):

Find the x and y components of the coyote’s average acceleration during that time.

OpenStudy (anonymous):

Let north be the y axis and east be the x axis. starting velocity v1 = (8, 0) second velocity v2 =(0, -8.80) change in velocity vector = (0, -8.80) -(8, 0) time needed for change = 4.20 s acceleration vector = change in velocity vector / 4.20 s